A cyclist is riding at a speed v on the level road. As he reaches a sharp circular corner of

radius r, he maintains the same speed. If the coefficient of friction between the tires and

the road is µ

(a) Show that the maximum permissible speed of the cyclist is

𝒗 = √𝝁𝒓𝒈

(b) If the cyclist’s speed was 20km/h, radius of the circular corner 15m and coefficient of

friction 0.35, show that the cyclist will skid away.


1
Expert's answer
2021-05-26T10:31:07-0400

(a) As the cyclist makes the turn, the friction force provides necessary centripetal force:

"F_c=F_{fr},""\dfrac{mv^2}{r}=\mu mg,""v=\sqrt{\mu rg}."

(b) Let's find the maximum permissible speed of the cyclist:


"v=\sqrt{\mu rg}=\sqrt{0.35\cdot15\ m\cdot9.8\ \dfrac{m}{s^2}}=7.17\ \dfrac{m}{s}."

The cyclist’s speed as he reaches a sharp circular corner was "20\ \dfrac{km}{h}" or "5.5\ \dfrac{m}{s}." As we can see, "v_{cyclist}<v_{max}," therefore, the cyclist will not skid away.